# Drug Discovery SMILES-to-Pharmacokinetics Diffusion Models with Deep Molecular Understanding

### v1

@@ Abstract
Artificial intelligence (AI) is increasingly used in every stage of drug development. One challenge facing drug discovery AI is that drug pharmacokinetic (PK) datasets are often collected independently from each other, often with limited overlap, creating data overlap sparsity. Data sparsity makes data curation difficult for researchers looking to answer research questions in poly-pharmacy, drug combination research, and high-throughput screening. We propose Imagand, a novel SMILES-to-Pharmacokinetic (S2PK) diffusion model capable of generating an array of PK target properties conditioned on SMILES inputs. We show that Imagand-generated synthetic PK data closely resembles real data univariate and bivariate distributions, and improves performance for downstream tasks. Imagand is a promising solution for data overlap sparsity and allows researchers to efficiently generate ligand PK data for drug discovery research. Code is available at https://github.com/bing1100/Imagand.
@@ 1 Introduction
Generative AI is set to transform drug discovery, where it may cost \$2-3 billion dollars and 10-15 years to bring a single drug candidate to market [30]. Generative AI for high-throughput screening (HTS) of ligand candidates reduces drug development costs and is changing how ligands are designed and tested [41]. Initial success of drug discovery AI has been in drug repurposing [50, 36], drug-target interaction [33], drug response prediction [40], poly-pharmacy [61], and the generation of synthetic ligands and drug properties [52, 25]. Thus far, what has advanced drug discovery AI is a continued effort towards open data for training and testing [27, 8, 17].
Data collection for drug discovery through assay panels is expensive and time-consuming. Although there are clear advances toward standardization and dissemination of pre-clinical, clinical, and chemical datasets [31, 27, 37], challenges arise when merging and linking these datasets together [46]. Collected independently, drug discovery datasets often have limited overlap, which poses a challenge for researchers looking to answer research questions requiring data from multiple datasets. One notable example is the study of drug combinations and poly-pharmacy [46].
Recent advances in drug discovery AI have utilized Denoising Diffusion Probabilistic Models (DDPMs) [21], which yield a new class of diffusion models capable of generating ligand structures [18, 52, 58, 28]. Hu et al. [25] have shown that diffusion models can generate pharmacokinetic (PK) properties alongside the ligand diffusion pipeline with promising results. Multi-modal diffusion models such as Text-to-Image and Text-to-Video diffusion models have been demonstrated to generate high-quality photorealistic data [44, 23]. Azizi et al. [4] have also shown that synthetic data from diffusion models have improved performance in ImageNet classification tasks. Inspired by these advances, we propose Imagand, which can generate an array of 12 PK target properties from 10 PK datasets conditioned on learned SMILES embeddings. Specifically, our contributions are as follows:
@@
  • We develop a noise model that creates a prior distribution closer to the true data distribution, which makes training easier [52].
  • We show that synthetic data generated from our Imagand model closely resembles the univariate and bivariate distributions of the real data and improves performance on downstream tasks.
@@ 2 Background
Diffusion methods use families of probability distributions to model complex datasets for computationally tractable learning, sampling, inference and evaluation [18]. Denoising Diffusion Probabilistic Models (DDPM) [21] first systematically destroy the structure in the data through a forward process, and then in a reverse process, learn how to restore the structure in the data from noise. Recent literature has covered many advances in small-molecule generation using diffusion models [26, 24, 45, 52].
DDPMs can be combined with graph networks. Conditional Diffusion models are based on discrete Graph Structures (CDGS) and can be used to generate small-molecule graphs with similar distribution to real small-molecules [26]. Digress [52] combines graph transformers [16] and discrete diffusion [3] for molecular-conditioned small-molecule generation. Digress utilizes graph-level properties such as cycles, and spectral and molecular features to augment the input, improving training and sampling performance. Conditional generation has shown benefits for diffusion Text-to-Image and Text-to-Video models [44, 23]. In Text-to-Image and Text-to-Video models, learned embeddings from Large Language Models (LLMs) are utilized as input to lend deep language understanding to the diffusion models.
PK broadly describes what the body does to a drug regarding absorption (how the body absorbs the drug), bioavailability (the extent the active drug enters circulation), distribution (how the drug distributes in tissue), metabolism (how the body breaks down the drug), and excretion (how the drug is removed from the body). As issues related to PK properties are the primary drivers for compound attrition for small-molecule drug development [32], accurate PK computational tools are critical and have advanced in recent times [55, 11, 2]. Physiologically-based pharmacokinetics (PBPK) offers the modelling of PK properties using mathematical equations representing the human body [43]. PBPK rely on expensive in-vitro and in-vivo human and animal experiments and cannot be utilized in high-throughput screening across large numbers of ligands (10K to 100K drugs per day) [39].
Extending many PK properties across large arrays of ligands can be costly given the expense associated with data collection for drugs. Consequently, oftentimes only small sets of ligands can be feasibly tested for target property data collection studies, leading to minimal overlap between collected datasets [46]. This challenge poses barriers for scientists interested in answering research questions requiring data across multiple datasets, such as in poly-pharmacy and drug combination research.
@@ 3 Methodology
Imagand is a S2PK diffusion model conditioned on learned SMILES embeddings from SMILES encoder models to generate target PK properties. We utilize a typical diffusion process as formulated in Ho et al. [21] with only modifications to the choice of noise model to better capture distribution priors. In the following subsections, we describe each of these components in detail.
@@ 3.1 Pre-trained SMILES Encoder
S2PK diffusion models need powerful semantic SMILE encoders to capture the complexity of arbitrary chemical structure inputs. Given the sparsity and small size of PK datasets, encoders trained on specific SMILES-Pharmacokinetic pairs are infeasible [27]. Many transformer-based foundational models such as ChemBERTa [9, 1], SMILES-BERT [54], and MOLGPT [5] have been pre-trained to deeply understand molecular and chemical structures and properties. After pre-training, these foundational models can then be fine-tuned for various downstream molecular tasks. Language models trained on SMILES-only corpus, significantly larger than SMILES-Pharmacokinetic data, learn a richer and wider distribution of molecular and chemical structures.
We test SMILES embeddings from ChemBERTa [1], T5 [42], and DeBERTa [19] trained on SMILES-only corpora. We further test and compare embedding performance for SMILES embedding from ChemBERTa trained either on ZINC (100K molecules) [29] or PubChem (10M molecules) [31] SMILES corpora. All SMILES embedding models were collected through the Huggingface [57] Model Hub. As ChemBERTa, T5, and DeBERTa are all trained on a wide array of SMILES, embeddings from these models are an effective way to inject deep molecular understanding into our diffusion model. Similar to Saharia et al. [44], we freeze the weights of our embedding models. Because embeddings are computed offline, freezing the weights minimizes computation and memory footprint for embeddings during model training.
3.2 Diffusion Model
3.2.1 Base Model
Imagand resembles a typical vision transformer architecture [14]; see Figure 1. 1D patches are computed from the classifier-free guidance of SMILES embeddings and concatenated with PK class tokens. Diffusion step embeddings are generated using sinusoidal position encodings [51]. Patches are then fed alongside sinusoidal step embeddings [22] to a transformer base. We mask out missing values when computing the loss for the model only to flow gradients and learn from non-missing PK values during training. Exponential Moving Average (EMA) [49] is applied to the base model during training to generate the final model used for sampling.
@@ 3.2.2 Classifier-free Guidance
Classifier guidance uses gradients from a pre-trained model to improve quality while reducing diversity in conditional diffusion models during sampling [12]. Classifier-free guidance [20] is an alternative technique that avoids this pre-trained model by jointly training a diffusion model on conditional and unconditional objectives via dropping the condition (i.e. with 10% probability). We condition all diffusion models on learned SMILES embedding and sinusoidal time embeddings using classifier-free guidance through dropout [20, 48].
3.2.3 Static Thresholding
We apply elementwise clipping the PK predictions to $[-1,1]$ as static thresholding, similar to Saharia et al. [44], Ho et al. [21]. Since PK data is min-max scaled to the same $[-1,1]$ range as a preprocessing step, static thresholding is essential to prevent the generation of invalid and out-of-range PK values.
3.3 Discrete Local Gaussian Noise Model
The choice of noise models may have a substantial impact on performance; using a prior distribution close to the true data distribution can make training easier [52]. As PK properties do not always follow a Gaussian or uniform noise model, we propose a noise model called Discrete Local Gaussian Noise (DLGN). DLGN decomposes complex PK distributions as Gaussian distributions within discrete bins. Discrete bins are modelled as discrete random variables.
3.3.1 Discrete Bin Random Variables
Given a PK distribution, we first apply data binning to produce $N$ bins. We define a discrete probability density function (PDF) on the bins, given in Equation 1, where $X_{i}$ represents the set of PK values in bin $i$, where $0\leq i\leq N$.
$P(X_{i})=\frac{|X_{i}|}{\sum^{N}_{j=0}{|X_{j}|}}$ (1)
3.3.2 Local Gaussian Distribution
For each $X_{i}$ bin, we define a Gaussian distribution $\hat{X}_{i}$ as denoted in Equation 2 with mean and standard deviation computed from $X_{i}$ bin.
$\hat{X}_{i}\sim\mathcal{N}(\mu(X_{i}),\sigma^{2}(X_{i}))$ (2)
Once a bin $X_{j}$ is chosen randomly from the discrete PDF (Equation 1), we then sample and return noise from the local bin Gaussian distribution $\hat{X}_{j}$. DLGN noise better resembles the prior distribution compared to Gaussian or uniform noise models, so training the model is faster.
@@ 3.4 Pharmacokinetic Datasets
All PK datasets are collected from TDCommons [27]. We select PK datasets suitable for regression from the absorption, distribution, metabolism, and excretion (ADME) and Toxicity categories.
@@
Caco-2 [53] is an absorption dataset containing rates of 906 drugs passing through the Caco-2 cells, approximating the rate at which the drugs permeate through the human intestinal tissue. Lipophilicity [59] is an absorption dataset that measures the ability of 4,200 drugs to dissolve in a lipid (e.g. fats, oils) environment. AqSolDB [47] is an absorption dataset that measures the ability of 9,982 drugs to dissolve in water. FreeSolv [35] is an absorption dataset that measures the experimental and calculated hydration-free energy of 642 drugs in water.
@@
Plasma Protein Binding Rate (PPBR) [56] is a distribution dataset of percentages for 1,614 drugs on how they bind to plasma proteins in the blood. Volume of Distribution at steady state (VDss) [34] is a distribution dataset that measures the degree for 1,130 drugs on their concentration in body tissue compared to their concentration in blood.
@@
Half Life [38] is an excretion dataset for 667 drugs on the duration for the concentration of the drug in the body to be reduced by half. Clearance [13] is an excretion dataset for around 1,050 drugs on two clearance experiment types, microsome and hepatocyte. Drug clearance is defined as the volume of plasma cleared of a drug over a specified time [27].
@@
Acute Toxicity (LD50) [60] is a toxicity dataset that measures the most conservative dose for 7,385 drugs that can lead to lethal adverse effects. hERG Central [15] is a toxicity dataset that measure the blocking of Human ether-à-go-go related gene (hERG) for 306,893 drugs. hERG is crucial for the coordination of the heart’s beating. hERG contains percentages inhibitions at $1\mu{}M$ and $10\mu{}M$.
Trainset-Testset
Data Metric Syn-Syn Syn-Real Real-Real
Caco2 MSE 0.158 0.131 0.634
R2 -0.069 0.137 -3.215
PCC 0.264 0.426 0.352
Lipo. MSE 0.156 0.150 0.167
R2 0.091 0.138 0.04
PCC 0.368 0.409 0.499
AqSol MSE 0.084 0.080 0.075
R2 0.507 0.533 0.564
PCC 0.713 0.731 0.756
FSolv MSE 0.194 0.165 0.624
R2 -0.139 0.078 -2.501
PCC 0.207 0.391 0.383
PPBR MSE 0.270 0.263 3.527
R2 -0.143 -0.06 -13.31
PCC 0.162 0.223 0.095
VDss MSE 0.201 0.196 0.535
R2 -0.132 -0.015 -1.771
PCC 0.206 0.298 0.234
Trainset-Testset
Data Metric Syn-Syn Syn-Real Real-Real
Half MSE 0.248 0.261 0.525
R2 -0.320 -0.275 -1.589
PCC 0.039 0.034 0.156
Cl. (H) MSE 0.437 0.433 1.863
R2 -0.279 -0.200 -4.24
PCC 0.053 0.096 0.109
Cl. (M) MSE 0.207 0.209 0.717
R2 -0.155 -0.043 -2.599
PCC 0.173 0.253 0.132
LD50 MSE 0.102 0.100 0.105
R2 0.253 0.277 0.240
PCC 0.517 0.537 0.542
hRG.1 MSE 0.126 0.127 0.136
R2 -0.111 -0.108 -0.189
PCC 0.060 0.062 0.062
hRG.10 MSE 0.115 0.115 0.121
R2 -0.030 -0.023 -0.081
PCC 0.189 0.196 0.212
Table 1: Comparing drug discovery regression performances between different combinations of augmented and real train sets, and augmented and real test sets. Values are averaged over 30 trials with the best scores on the real test set bolded. Statistical significance (underlined) is computed with a two-tailed T-test. Hyphenated pairs correspond to the train and test set used in the experiment.

### v2

@@ Abstract
The role of Artificial Intelligence (AI) is growing in every stage of drug development. Nevertheless, a major challenge in drug discovery AI remains: Drug pharmacokinetic (PK) datasets collected in different studies often exhibit limited overlap, creating data overlap sparsity. Thus, data curation becomes difficult, negatively impacting downstream research investigations in high-throughput screening, polypharmacy, and drug combination. We propose Imagand, a novel SMILES-to-Pharmacokinetic (S2PK) diffusion model capable of generating an array of PK target properties conditioned on SMILES inputs that exhibit data overlap sparsity. We show that Imagand-generated synthetic PK data closely resembles real data univariate and bivariate distributions, and can adequately fill in gaps among PK datasets. As such, Imagand is a promising solution for data overlap sparsity and may improve performance for downstream drug discovery research tasks. Code available at: https://github.com/GenerativeDrugDiscovery/imagand
Machine Learning, ICML
@@ 1 Introduction
Generative AI is set to transform drug discovery, where it may cost \$2-3 billion dollars and 10-15 years to bring a single drug candidate to market (Kim et al., 2021). Generative AI for high-throughput screening (HTS) of ligand candidates reduces drug development costs and is changing how ligands are designed and tested (Pushpakom et al., 2019). Initial success of drug discovery AI has been in drug repurposing (Thafar et al., 2022; Morselli Gysi et al., 2021), drug-target interaction (Lian et al., 2021), drug response prediction (Pouryahya et al., 2022), poly-pharmacy (Žitnik et al., 2015), and the generation of synthetic ligands and drug properties (Vignac et al., 2023; Hu et al., 2024). Thus far, what has advanced drug discovery AI is a continued effort towards open data for training and testing (Huang et al., 2021; Brown et al., 2019; Gaulton et al., 2017).
Data collection for drug discovery through assay panels is expensive and time-consuming. Although there are clear advances toward standardization and dissemination of pre-clinical, clinical, and chemical datasets (Kim et al., 2023; Huang et al., 2021; Nusinow et al., 2020), challenges arise when merging and linking these datasets together (Scoarta et al., 2023). Collected independently, drug discovery datasets often have limited overlap, which poses a challenge for researchers looking to answer research questions requiring data from multiple datasets. One notable example is the study of drug combinations and polypharmacy (Scoarta et al., 2023).
Recent advances in drug discovery AI have utilized Denoising Diffusion Probabilistic Models (DDPMs) (Ho et al., 2020), which yield a new class of diffusion models capable of generating ligand structures (Guo et al., 2023; Vignac et al., 2023; Wu et al., 2022; Igashov et al., 2022). Hu et al. (2024) have shown that diffusion models can generate pharmacokinetic (PK) properties alongside the ligand diffusion pipeline with promising results. Inspired by these advances, we propose Imagand, which can generate 12 PK target properties from 10 PK datasets conditioned on learned SMILES embeddings. Specifically, our contributions are as follows:
@@
  • We develop a noise model that creates a prior distribution closer to the true data distribution, improving performance.
  • We show that synthetic data generated from our Imagand model has univariate and bivariate distributions closely matching the real data and improves machine learning efficiency.
@@ 2 Background
Diffusion methods use families of probability distributions to model complex datasets for computationally tractable learning, sampling, inference and evaluation (Guo et al., 2023). Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020) first systematically destroy the structure in the data through a forward process, and then in a reverse process, learn how to restore the structure in the data from noise. Recent literature has covered many advances in small-molecule generation using diffusion models (Huang et al., 2023; Hoogeboom et al., 2022; Satorras et al., 2021; Vignac et al., 2023).
PK broadly describes what the body does to a drug regarding absorption (how the body absorbs the drug), bioavailability (the extent the active drug enters circulation), distribution (how the drug distributes in tissue), metabolism (how the body breaks down the drug), and excretion (how the drug is removed from the body). As issues related to PK properties are the primary drivers for compound attrition for small-molecule drug development (Kola, 2008), accurate PK computational tools are critical and have advanced in recent times (Waring et al., 2015; Davies et al., 2020; Ahmed et al., 2021). Physiologically-based pharmacokinetics (PBPK) offers the modelling of PK properties using mathematical equations representing the human body (Sager et al., 2015). PBPK rely on expensive in-vitro and in-vivo human and animal experiments and cannot be utilized in high-throughput screening across large numbers of ligands (10K to 100K drugs per day) (Obrezanova, 2023).
Extending many PK properties across large arrays of ligands can be costly given the expense associated with data collection for drugs. Consequently, oftentimes only small sets of ligands can be feasibly tested for target property data collection studies, leading to minimal overlap between collected datasets (Scoarta et al., 2023). Comparing the 11 PK datasets we use in this study, Table 1 shows the minimal overlap sparsity between all of the datasets. This challenge poses barriers for scientists interested in answering research questions requiring data across multiple datasets, such as in poly-pharmacy and drug combination research.
@@ 3 Methodology
The choice of noise models for noising in diffusion models may have a substantial impact on performance; using a prior distribution close to the true data distribution can make training easier (Vignac et al., 2023). As PK properties do not always follow a Gaussian or uniform noise model, we propose a noise model called Discrete Local Gaussian Noise (DLGN).
3.1 Discrete Local Gaussian Sampling
Discrete local gaussian (DLG) sampling is based on principles of inverse transform sampling. Inverse transform sampling is a method for sampling from any probability distribution given its cumulative distribution function. For any variable $X\in\mathbb{R}$, the random variable $F_{X}^{-1}(U)$ has the same distribution as $X$, where $F_{X}^{-1}$ is the generalized inverse of the cumulative distribution function $F_{X}$ of $X$ and $U\sim Unif[0,1]$.
As the true distribution is not always available, Discrete Local Gaussian Sampling $\mathbb{DLGS}$ looks to discretely approximate the cumulative distribution function by combining binning and Gaussian noise given by Theorem 3.1 (with proof Appendix A). Given $N$ bins, a discrete cumulative distribution function $\hat{F}_{X_{N}}^{-1}$ can be constructed. With $\sigma$ as a scaling factor, we can then define DLG sampling $\phi(X=x)$ as:
$\mathbb{\phi}(X=x)$ $:=\mathcal{N}(x;\hat{F}_{X_{N}}^{-1}(U),\frac{\sigma}{N^{2}}\mathbf{I})$ (1)
Theorem 3.1.
Let $D_{KL}(P||Q)$ be the Kullback-Leibler (KL) divergence for a model probability distribution $Q$ and the true probability distribution $P$. For any distribution $X$, we then have:
$lim_{N\to\infty}\ D_{KL}(X\ ||\ \mathbb{\phi}(X=x))=0$
3.2 Imagand Model
Imagand is an S2PK diffusion model conditioned on learned SMILES embeddings from SMILES encoder models to generate target PK properties. Imagand improves accuracy by training a diffusion process (Ho et al., 2020) utilizing custom noise models.
3.2.1 Diffusion Model
Given samples from a data distribution $q(x_{0})$, we are interested in learning a model distribution $p_{\theta}(x_{0})$ that approximates $q(x_{0})$ and is easy to sample from. Ho et al. (2020) considers the following Markov chain with Gaussian transitions parameterized by a decreasing sequence $\alpha_{1:T}\in(0,1]^{T}$:
$q(x_{1:T}|x_{0}):=\mathcal{N}(x_{1:T}|\sqrt{\alpha_{1:T}}x_{0},(1-\alpha_{1:T})\mathbf{I})$ (2)
This is called the forward process, whereas the latent variable model $p_{\theta}(x_{0:T})$ is the generative process, approximating the reverse process $q(x_{t-1}|x_{t})$. The forward process of $x_{t}$ can be expressed as a linear combination of $x_{0}$ and noise variable $\epsilon$:
$x_{t}=\sqrt{\alpha_{t}}x_{0}\ +\ \sqrt{1-\alpha_{t}}\epsilon$ (3)
We experiment with noise models $\epsilon$, testing with Average, Uniform, Gaussian, and DLG sampling approaches. Empirically, from our ablation studies, we find that DLG sampling improves performance over other noise models, as it better approximates the original data. We train with the simplified objective:
$L(\epsilon_{\theta}):=\sum_{t=1}^{T}\mathbb{E}_{x_{0}\sim q(x_{0}),\epsilon_{t}}[||\epsilon_{\theta}^{(t)}(x_{t})-\epsilon_{t}||_{2}^{2}]$ (4)
Where $\epsilon_{\theta}:=\{\epsilon_{\theta}^{(t)}\}_{t=1}^{T}$ is a set of T functions, indexed by t, each with trainable parameters $\theta^{(t)}$. Convergence analysis of utilizing DLG sampling as a noise model is provided in Appendix C.
3.2.2 Architecture
Imagand resembles a typical vision transformer architecture (Dosovitskiy et al., 2021); see Figure 1. 1D patches are computed from the classifier-free guidance of SMILES embeddings and concatenated with PK class tokens. Diffusion step embeddings are generated using sinusoidal position encodings (Vaswani et al., 2023). Patches are then fed alongside sinusoidal step embeddings (Ho et al., 2021) to a transformer base. We mask out missing values when computing the loss for the model only to flow gradients and learn from non-missing PK values during training. Exponential Moving Average (EMA) (Tarvainen & Valpola, 2018) is applied to the base model during training to generate the final model used for sampling.
@@ 3.3 Pre-trained SMILES Encoder
S2PK diffusion models need powerful semantic SMILE encoders to capture the complexity of arbitrary chemical structure inputs. Given the sparsity and small size of PK datasets, encoders trained on specific SMILES-Pharmacokinetic pairs are infeasible (Huang et al., 2021). Many transformer-based foundational models such as ChemBERTa (Chithrananda et al., 2020; Ahmad et al., 2022), SMILES-BERT (Wang et al., 2019), and MOLGPT (Bagal et al., 2021) have been pre-trained to deeply understand molecular and chemical structures and properties. After pre-training, these foundational models can then be fine-tuned for various downstream molecular tasks. Language models trained on SMILES-only corpus, significantly larger than SMILES-Pharmacokinetic data, learn a richer and wider distribution of molecular and chemical structures.
We test SMILES embeddings from ChemBERTa (Ahmad et al., 2022), T5 (Raffel et al., 2023), and DeBERTa (He et al., 2021) trained on SMILES-only corpora. We further test and compare embedding performance for SMILES embedding from ChemBERTa trained either on ZINC (100K molecules) (Irwin & Shoichet, 2005) or PubChem (10M molecules) (Kim et al., 2023) SMILES corpora. All SMILES embedding models were collected through the Huggingface (Wolf et al., 2020) Model Hub. As ChemBERTa, T5, and DeBERTa are all trained on a wide array of SMILES, embeddings from these models are an effective way to inject deep molecular understanding into our diffusion model. Similar to Saharia et al. (2022), we freeze the weights of our embedding models. Because embeddings are computed offline, freezing the weights minimizes computation and memory footprint for embeddings during model training.
@@ 3.3.1 Classifier-free Guidance
Classifier guidance uses gradients from a pre-trained model to improve quality while reducing diversity in conditional diffusion models during sampling (Dhariwal & Nichol, 2021). Classifier-free guidance (Ho & Salimans, 2022) is an alternative technique that avoids this pre-trained model by jointly training a diffusion model on conditional and unconditional objectives via dropping the condition (i.e. with 10% probability). We condition all diffusion models on learned SMILES embedding and sinusoidal time embeddings using classifier-free guidance through dropout (Ho & Salimans, 2022; Srivastava et al., 2014).
Data Overlap Cardinality 1 2 3 4 5 6 7 8 9 10 11
Number of Data Points 322235 3598 404 1110 34 105 27 4 14 2 0
Table 1: Number of data points for each data overlap cardinality over 11 PK datasets. Data overlap cardinality represents the number of datasets a data point is in. The number of data points in data overlap cardinality 1 represents the number of data points only in one of the 11 PK datasets. As data overlap cardinality increases, the number of data points greatly drops, describing the data overlap sparsity phenomenon. This lack of data negatively impacts downstream research investigations in high-throughput screening, poly-pharmacy, and drug combination.
4 Experiments
In the following, we describe the model training details and compare our synthetic data to real data, in terms of machine learning efficiency (MLE) and univariate and bivariate statistical distributions. We then discuss ablation studies and key findings. The metrics for MLE, univariate, and bivariate evaluations are further defined in their respective subsections.
We compare Imagand with baselines of Conditional GAN (cGAN) (Mirza & Osindero, 2014) and Syngand (Hu et al., 2024). Similar to in Imagand, SMILES-embeddings from a pre-trained T5 model are used conditionally by the cGAN model to generate PK properties as output for a specific drug. Additional training details and baseline results are provided in Appendix B and Appendix D.
@@ 4.1 Pharmacokinetic Datasets
All 11 PK datasets are collected from TDCommons (Huang et al., 2021). We select PK datasets suitable for regression from the absorption, distribution, metabolism, and excretion (ADME) and Toxicity categories. Looking over 11 PK datasets for target property screening, Table 1 shows data overlap sparsity as the data overlap cardinality increases, the number of data points greatly drops.
@@
Caco-2 (Wang et al., 2016) is an absorption dataset containing rates of 906 drugs passing through the Caco-2 cells, approximating the rate at which the drugs permeate through the human intestinal tissue. Lipophilicity (Wu et al., 2018) is an absorption dataset that measures the ability of 4,200 drugs to dissolve in a lipid (e.g. fats, oils) environment. AqSolDB (Sorkun et al., 2019) is an absorption dataset that measures the ability of 9,982 drugs to dissolve in water. FreeSolv (Mobley & Guthrie, 2014) is an absorption dataset that measures the experimental and calculated hydration-free energy of 642 drugs in water.
@@
Plasma Protein Binding Rate (PPBR) (Wenlock & Tomkinson, 2016) is a distribution dataset of percentages for 1,614 drugs on how they bind to plasma proteins in the blood. Volume of Distribution at steady state (VDss) (Lombardo & Jing, 2016) is a distribution dataset that measures the degree for 1,130 drugs on their concentration in body tissue compared to their concentration in blood.
@@
Half Life (Obach et al., 2008) is an excretion dataset for 667 drugs on the duration for the concentration of the drug in the body to be reduced by half. Clearance (Di et al., 2012) is an excretion dataset for around 1,050 drugs on two clearance experiment types, microsome and hepatocyte. Drug clearance is defined as the volume of plasma cleared of a drug over a specified time (Huang et al., 2021).
@@
Acute Toxicity (LD50) (Zhu et al., 2009) is a toxicity dataset that measures the most conservative dose for 7,385 drugs that can lead to lethal adverse effects. hERG Central (Du et al., 2011) is a toxicity dataset that measure the blocking of Human ether-à-go-go related gene (hERG) for 306,893 drugs. hERG is crucial for the coordination of the heart’s beating. hERG contains percentages inhibitions at $1\mu{}M$ and $10\mu{}M$.
Model
Data Metric cGAN Sygd* Imgd Real
Caco2 MSE 0.165 0.276 0.131 0.634
R2 -0.08 -3.35 0.137 -3.215
PCC 0.338 0.302 0.426 0.352
Lipo. MSE 0.141 0.313 0.150 0.167
R2 0.194 0.126 0.138 0.04
PCC 0.469 0.181 0.409 0.499
AqSol MSE 0.074 0.107 0.08 0.075
R2 0.565 0.348 0.533 0.564
PCC 0.755 0.681 0.731 0.756
FSolv MSE 0.198 0.182 0.165 0.624
R2 -0.09 -0.023 0.078 -2.501
PCC 0.422 0.515 0.391 0.383
PPBR MSE 0.26 0.361 0.263 3.527
R2 -0.082 -1.36 -0.06 -13.31
PCC 0.225 0.125 0.223 0.095
VDss MSE 0.209 0.307 0.196 0.535
R2 -0.064 -0.843 -0.015 -1.771
PCC 0.306 0.189 0.298 0.234
Model
Data Metric cGAN Sygd* Imgd Real
Half MSE 0.284 0.437 0.261 0.525
R2 -0.536 -0.831 -0.275 -1.589
PCC 0.134 0.065 0.034 0.156
Cl.(H) MSE 0.431 0.563 0.433 1.863
R2 -0.153 -1.14 -0.200 -4.24
PCC 0.144 0.032 0.096 0.109
Cl.(M) MSE 0.203 0.278 0.209 0.717
R2 -0.037 -1.71 -0.043 -2.599
PCC 0.25 0.189 0.253 0.132
LD50 MSE 0.103 0.111 0.100 0.105
R2 0.252 0.298 0.277 0.240
PCC 0.526 0.558 0.537 0.542
hRG.1 MSE 0.132 0.121 0.127 0.136
R2 -0.135 -0.55 -0.108 -0.189
PCC 0.035 0.060 0.062 0.062
hRG.10 MSE 0.134 0.106 0.115 0.121
R2 -0.182 -0.075 -0.023 -0.081
PCC 0.207 0.200 0.196 0.212
Table 2: Comparing drug discovery Machine Learning Efficiency (MLE) regression performances between different models and with real train data. Mean Squared Error (MSE), R-Squared (R2), and Pearson Correlation Coefficient (PCC) values are averaged over 30 trials, with the best scores on the real test set bolded. *Syngand R2 and PCC results are scale-adjusted relative to Real-Real with cGAN and Imagand results.